OpenStudy (anonymous):

Solve the equation. 1/4-t/2=15/8

OpenStudy (pradipgr817):

-13/4

OpenStudy (anonymous):

1/4 - t/2 = 15/8 1/4 = 15/8 + t/2 1/4 - 15/8 = t/2 2/8 - 15/8 = t/2 -13/8 = t/2 -13/4 = t